OpenStudy (anonymous):

A cylindrical can has a horizontal base of radius 3.4 cm. It contains sufficient water so that when a sphere is placed inside, the water just covers the sphere. If the sphere fits exactly into the can, calculate (a) The total surface area of the can in contact with the water when the shpere is inside, (b) The depth of the water in the can before the sphere was put inside.

OpenStudy (anonymous):



OpenStudy (anonymous):

this?

OpenStudy (anonymous):

There's no diagram with it.

OpenStudy (anonymous):

question no.22

OpenStudy (baru):



OpenStudy (baru):

if the sphere fits into the can exactly, then diameter of the can= diameter of the sphere

OpenStudy (anonymous):

I think the r of sphere is 6.8, h of cylindrical is 6.8 but I'm russian and can not understand this exactly

OpenStudy (baru):

for part (b) it is just volume of water= volume of cylinder - volume of sphere

OpenStudy (baru):

and yes, height also equal to diameter :)

OpenStudy (anonymous):

cylindrical S = h*pi*r + 2*pi*r^2

OpenStudy (baru):

i'm not sure about part (a)

OpenStudy (anonymous):

WAIT S = 2 * pi *r*h + 2 * pi * r^2 But it's can so S = 2*pi*r*h + pi*r^2= pi*r*(2h+r)

OpenStudy (anonymous):

S= 2*3.142*3.4*6.8+3.142*3.4^2 =145.29+36.32 =181.61 ^Correct answer.

OpenStudy (baru):

oh ok, so its just the surface area of the entire can

OpenStudy (baru):

@Palvasha are we done?

OpenStudy (anonymous):

Yeah. Thank you both of you.

OpenStudy (baru):

:)